\begin{thebibliography}{99}

\bibitem{reference1}
N. Cabibbo, Unitary symmetry and leptonic decays, Phys. Rev. Lett. 10, 531, 1963.

\bibitem{reference2}
S. Navas et al, Particle Data Group: $CKM$ quark-mixing matrix. Phys. Rev. D 110 030001, 2024.

\bibitem{reference3}
P. Denton and R. Pestes, The impact of different parameterizations on the interpretation of $CP$ violation in neutrino oscillations, JHEP 05 2021 039.

\bibitem{reference4}
A. Hocker, H. Lacker, S. Palace and F. Diberder, A new approach to a global fit of the CKM matrix. Eur. Phys. J. C 21, 225-229 2001.

%\bibitem{reference5}
%H. Ke, T. Liu and X. Li, Determination of the mixing between active neutrinos and sterile neutrino through the quark-lepton complementarity and self-complementarity, Phys. Rev. D 90, 053009, 2014.

\bibitem{reference6}
D. Atwood, S. Gupta and A. Soni, Detecting fourth generation quark in hadron colliders.  JHEP 1206 105, 2012.

\bibitem{reference7}
%A. Esposito, Rafael Krichevsky and A. Nicolis, The gravitational mass carried by sound waves, Phys. Rev. Lett. 122 084501 2019.
G. Kaur, A. Bagai, G. Ahuja and M. Gupta, Revisiting rep -resentations of quark mixing matrix.  Accepted for publication in Progress in Theoretical and Experimental Physics.
%I. Aitchison and A. Hey, Gauge theory in particle physics, Fourth edition, CRC Press, 2013.

\bibitem{reference8}
Y. Ahn, H. Cheng and S. Oh, Wolfenstein parameterization at higher order: seeming discrepancies and their resolution.  Phys. Lett. B. 703 571-575 2011. 

\bibitem{reference9}
M. Bona et al, Model-independent constraints on $\Delta F$=2 operators and the scale of new physics.  JHEP 0803 2008 049.

\bibitem{reference10}
M. Kobayashi, Exactly unitary Wolfenstein parameterization of flavor mixing.  Progr of Theor Phys 92 287-288, 1994.
%\bibitem{reference16}

\end{thebibliography}

\documentclass[%
 reprint,
 amsmath,amssymb,
 aps,
]{revtex4-2}

\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}

\begin{document}

\preprint{APS/123-QED}

\title{A constraint to the uncertainty associated with the elements in the quark mixing matrix}

\author{\small Jae Jun Kim}
 
\affiliation{
 Research and Data Analysis\\
 South Carolina Department of Education\\
 Columbia, South Carolina\\
The United States
 \\
}%


\date{\today}

\begin{abstract}
We present an extra condition by which the size of the uncertainty associated with the diagonal elements in the quark mixing matrix can be further constrained when the matrix is parameterized by a combination of three mixing angles and a Dirac phase.  Then we discuss how the correlations among the uncertainties of the diagonal elements can be used when we address issues such as the tension in the first row and the column of the matrix and discuss its implications.
\end{abstract}

\maketitle


\noindent
Since the formulation of quark mixing by Cabibbo, numerous studies have been conducted to better understand the mechanisms of the mixing \cite{reference1}.  In our study, as a continuation of the efforts, we present an extra constraint, by which the size of the uncertainty associated with the elements, the diagonal elements in particular, of the mixing matrix, can be correlated, taking into account the unitarity of the mixing matrix in a standard parameterization scheme, a scheme that is based on the isospin doublet.  Then we discuss how such correlations among the size of the uncertainties can be used when addressing issues such as the tension in the first row and the column of the matrix in the quark sector.


Mathematically, we can parameterize a unitary matrix in many different ways.  Among them, there are two well-known schemes of parameterizing the matrix: Wolfenstein parameterization, which was introduced as an approximate scheme first but made to be exact later and widely adopted when testing the unitarity \cite{reference10}, and the standard parameterizations illustrated in $PDG$ \cite{reference2}, where the matrix is constructed by a combination of $2 \times 2$ rotational matrices.  The two types of parameterization schemes have been widely adopted among researchers, given their suitability to our measuring the size of the relevant parameters and describing the mixing.  

Understanding and imposing more stringent limits on the unitarity of the mixing matrix has been considered as one such important topic \cite{reference2}.  When we trying to do so, asides from the correlation among parameters in various phase spaces, the following two conditions have been considered: The square sum of the elements in each column and row adds up to unity, and Jarlskog invariant \cite{reference3},
\begin{equation}
\begin{split}
J \equiv IM[V^{}_{11} \hspace{0.2mm} V^*_{12} \hspace{0.2mm} V^*_{21} \hspace{0.2mm} V^{}_{22}]
\end{split}
\end{equation}
, where $V_i$ is an element in the unitary matrix in Equation 4, needs to be held.  Equation 1 can be written as a function of four independent rotational parameters in the standard scheme of writing the matrix \cite{reference2}.  Imposing the two conditions along with that in different phase spaces has led us to constrain the size of the elements in the mixing matrix further when running fitters \cite{reference4}. 

Interestingly, most of the rules regarding the unitarity can be applied to the size of the central value of the elements of the matrix or that of the phase space associated with the central values.  Many studies focused on parameterizing the matrix that suits the phase space of the measurable quantities as functions of the parameters used when constructing the matrix.

That being said, there are not many studies conducted to explore conditions associated with the unitarity to that of the size of the uncertainty associated with the elements in the matrix, not the central value.  Can we possibly constrain the size of uncertainty, either in their leading-order or exact?  There, our study began.

We focus on exploring a possible correlation that exists among the size of the uncertainty associated with each element, not the central value, when parameterizing it in a combination of three $\theta$s and one $\delta$, the standard parameterization in \cite{reference2}, where unitarity is guaranteed.  For that, first we describe what we can do with the uncertainty part of the matrix using the two conditions mentioned above.  After that, we study the correlation among the size of the uncertainties and its relationship to some of the issues such as the tension in the first row and column of the matrix in the quark sector.  

When we parameterize the matrix, it is known that we need at least four parameters to be used, by reducing the number to half, demanding the unitarity, along with the conditions of rephrasing the phase.  With the minimum number of parameters, the entire matrix, including the size of uncertainty associated with the matrix elements, can be constructed as a function of four independent parameters and the uncertainty associated with them.  

Note that, when we describe mixing by a combination of three $2 \times 2$ rotational matrices under the standard parameterization scheme, the permutations among the parameters need to be considered.  In other words, the order by which $2 \times 2$ matrix comes matter when constructing the whole $3 \times 3$ mixing matrix \cite{reference3}.  For instance, when we construct the matrix by three $\theta$s and one $\delta$, we need to consider at least $6=3 \times 2 \times 1$ permutations.  As illustrated in \cite{reference3},
\begin{equation}
\begin{split}
&U = U_{1} \hspace{0.1mm} U_{2} \hspace{0.1mm} U_{3} = U_{1} \hspace{0.1mm} U_{3} \hspace{0.1mm} U_{2} = U_{2} \hspace{0.1mm} U_{1} \hspace{0.1mm} U_{3}\\
&= U_{2} \hspace{0.1mm} U_{3} \hspace{0.1mm} U_{1} = U_{3} \hspace{0.1mm} U_{1} \hspace{0.1mm} U_{2} = U_{3} \hspace{0.1mm} U_{2} \hspace{0.1mm} U_{1}
\end{split}
\end{equation}
, where $U_{1,2,3}$ are $2 \times 2$ rotational matrices, parameterized by three $\theta$s and one $\delta$, as in other literature.

In fact, there are nine different ways to parameterize the matrix within the scheme of constructing the matrix with the three angles and the one phase \cite{reference6}.  It is nine since we need to consider the presence of the Dirac phase in addition to the permutation among the rotational matrices.  This happened to be not an issue in parameterization schemes such as Wolfenstein parameterization though, since the parameterization is not based on rotations but based on scalings \cite{reference8}.

It is not too difficult to see that all six cases of the permutation return different sizes of three $\theta$ s and one $\delta$ but the same size of the elements in the unitary mixing matrix.  There, our question arises: Are there functions in which the size of the uncertainty, or the relative uncertainty in some combinations of the elements, can stay invariant, either in the leading order or exact, when the unitarity of the mixing matrix is assumed by the two main conditions mentioned in the beginning?

Interestingly, there is one such function that we could identify, not directly the size of the uncertainty, but related to the ratio of the elements, which can be utilized to constrain the size of the uncertainty among the elements further to some extent.  The one we explore in this work can begin with a combination of four sine functions.  Taking into account the expressions that represent the size of the matrix elements in \cite{reference3}, it can be written as
\begin{equation}
\begin{split}
S_i = sin \theta^i_{23} sin \theta^i_{12} sin \theta^i_{13} sin \delta^i_{13}
\end{split}
\end{equation}
, where $i$ runs from 1 to 6, representing different ordering of $U_{1,2,3}$.  Note that we only consider a Dirac phase as a component in the expression.

Writing a $3 \times 3$ unitarity matrix as,
\begin{equation}
V=
\begin{array}{c|ccccc|c}
&& V_{11} & V_{12} & V_{13} && \\
&& V_{21} & V_{22} & V_{23} && \\
&& V_{31} & V_{32} & V_{33} && \\
\end{array}
\end{equation}
, then it is not too difficult to see $S_i$ is not invariant, given that three $\theta$s and one $\delta$ can have different values when the rotational matrices are ordered differently.  

However, in all six cases, $S_i$ can be written as a function of a combination of Jarlskog invariant, along with the two out of three diagonal elements in the matrix as,
\begin{equation}
\begin{split}
&J = S_{1} \hspace{0.2mm} V_{1} \hspace{0.2mm} V_{3} = S_{2} \hspace{0.2mm} V_{1} \hspace{0.2mm} V_{2} = S_{3} \hspace{0.2mm} V_{2} \hspace{0.2mm} V_{3}\\
& = S_{4} \hspace{0.2mm} V_{1} \hspace{0.2mm} V_{2} = S_{5} \hspace{0.2mm} V_{2} \hspace{0.2mm} V_{3} = S_{6} \hspace{0.2mm} V_{1} \hspace{0.2mm} V_{3}
\end{split}
\end{equation}
, where the six cases represent the six different permutations of ordering the matrices in Equation 2.  Note that we use a single index for the diagonal elements,
\begin{equation}
V_{1} = V_{11}, V_{2} = V_{22}, V_{3} = V_{33} 
\end{equation}
.  Given the invariance of $J$, the size of the ratio of $S$ for the six different cases of permutation can be written as a ratio of the size of the two elements in the matrix.  For instance, taking the ratio of the first three cases of permutations, $S_{1} \hspace{0.2mm} V_{1} \hspace{0.2mm} V_{3}$, $S_{2} \hspace{0.2mm} V_{1} \hspace{0.2mm} V_{2}$ and $S_{3} \hspace{0.2mm} V_{2} \hspace{0.2mm} V_{3}$, we have,
\begin{equation}
\begin{split}
A_1 = \frac{S_1}{S_2} = \frac{V_3}{V_2}, \hspace{0.2mm}
A_2 = \frac{S_1}{S_3} = \frac{V_1}{V_2}, \hspace{0.2mm}
A_3 = \frac{S_2}{S_3} = \frac{V_1}{V_3}\\
\end{split}
\end{equation}
, where $J$ and one of $V_i$ cancel out from Equation 5, when taking the ratio.  The rest of the combination of $S_i$ can also be written as a function of one of the three $A_i$.  With that, we can define $C$, the relative size of the uncertainty, the size of uncertainty with respect to that of the central value of the ratio of $S$, as,
\begin{equation}
\begin{split}
C \sim \frac{1}{A_1} \hspace{0.2mm} \Delta A_1 \sim \frac{1}{A_2} \hspace{0.2mm} \Delta A_2 \sim \frac{1}{A_3} \hspace{0.2mm} \Delta A_3
\end{split}
\end{equation}
, which can be written as a function of cotangent.  Expanding the function and taking the leading order term for each $\theta$ and $\delta$, for all six cases of permutations, we end up with,
\begin{equation}
\begin{split}
\frac{1}{2}C \sim
\frac{1}{\theta_{23}}\Delta\hspace{0.3mm}\theta_{23} +
\frac{1}{\theta_{13}}\Delta\hspace{0.3mm}\theta_{13} +
\frac{1}{\theta_{12}}\Delta\hspace{0.3mm}\theta_{12} +
\frac{1}{\theta_{13}}\Delta\hspace{0.3mm}\delta_{13}
\end{split}
\end{equation}
, where $C$, the size of the uncertainty of the ratio of $S$, can stay invariant, in its leading order, no matter how we order the three $2 \times 2$ rotation matrix.  The size of the uncertainty due to not including the next leading order term in $C$ is $\sim$ $5\%$.  Note that we do not have index $i$ in $\theta$ and $\delta$ since the relative uncertainty for each parameter is going to be same.  

Having such leads us to an interesting implication: When we add up the size of the relative uncertainty of the two out of the three diagonal elements in the $3 \times 3$ unitarity mixing matrix, it stays invariant in the leading order, given the size of the relative uncertainty associated with each parameter, at least within the scheme of having a combination of three $\theta$s and one $\delta$, stays the same.

Equation 9 indicates that, given two out of three diagonal elements in the matrix, we can choose two different orderings in the parameterization in combination under which the size of the relative uncertainty of their ratio can be constrained.  This cannot be realized with a single diagonal element in the matrix.




In other words, when parameterizing the $3 \times 3$ matrix by the three rotational matrices, we cannot say for sure that the uncertainty size for a single diagonal element in the matrix is such in its leading order.  They need to be reported at least by two out of three diagonal elements, or at least the uncertainty on $V$ needs to be reported separately.

Having such patterns can be applied to some of the issues including the tension in the first row of the matrix in the quark sector.

The tension in essence is that the square sum of the size of the elements in the first row of the quark mixing matrix does not add up to unity with the uncertainty,
\begin{equation}
\begin{split}
|V_{1}|^2 + |V_{12}|^2 + |V_{13}|^2 = 0.9977 - 0.9991
\end{split}
\end{equation}
, where the first number is lower and the second number the upper limit.  The size of the uncertainty associated with each element is,
\begin{equation}
\begin{split}
&\Delta V_1 \sim 1.6 \times10^{-4}, \hspace{0.2mm} \Delta V_{12} \sim 6.8 \times10^{-4},\\
&\Delta V_{13} \sim 8.75 \times10^{-5}
\end{split}
\end{equation}
, as a fixed value returned by running a global fit such as that \cite{reference4}.  The central value and that of the uncertainty were measured in Wolfenstein parameterization.  

The question of whether the matrix is an unitary or not is beyond the scope of our study.  However, if some of the characteristics associated with the unitarity is assumed when running the global fits as that described in \cite{reference4} and \cite{reference9}, we may need additional constraints of their uncertainties being correlated as shown in Equation 9 needs to be taken into account.  The degree of impact of imposing the constraint as an extra condition on $\Delta V_1$ may depend on how it is correlated to the constraints in different phase spaces when running the fitter.

In summary, in addition to the common constraints that can be imposed when considering a matrix to be unitary, we may need to consider additional conditions such as the correlation between the size of the uncertainty associated with the diagonal elements in the mixing matrix.  It cannot be realized with a single diagonal element, but two out of three given the characteristic associated with the standard scheme of parameterizing the mixing matrix.  Such could be correlated with other constraints imposed when running a global fit, so it is rather difficult to say whether it is going to alleviate the tension or make it even worse, but it is worth trying to incorporate it and see how it works when running fitters.

As for future studies, one of them could be about expanding the work to the $4 \times 4$ matrix and to a higher generation, where we set the central value for the fourth generation to zero except for the diagonal element, calculate the uncertainty size introduced due to tensions in the framework of $3 \times 3$ and see if it returns a consistent result with the theoretical estimates \cite{reference6}.




\end{document}